OpenStudy (anonymous):

solve the equation f(x)=g(x) algebraically. f(x)= ln(x-4) g9x)=0

OpenStudy (anonymous):

so you need to solve: \[\huge 0=ln(x-4) \]

OpenStudy (anonymous):

can you write this in exponential form?

OpenStudy (anonymous):

how do you do that? :O

OpenStudy (anonymous):

any exponential equation can be written in logarithmic form and vice versa... \[\huge b^M=N\Leftrightarrow log_bN=M \]

OpenStudy (anonymous):

so in your case.. \[\huge ln(x-4)=0\Leftrightarrow e^0=x-4 \]

OpenStudy (anonymous):

do you know what \(\large e^0= \)????

OpenStudy (anonymous):

1? sorry late reply

OpenStudy (matheducatormcg):

ooops... I mean \[1=x-4\]

OpenStudy (lgbasallote):

\[e^0 = 1\] therefore \[1 = x - 4\] now solve for x